OpenStudy (issimplcalcus):

What would be the way you'd anti derive 25∫cos(arctan(1.2/x)?

OpenStudy (issimplcalcus):

I dont need a solution just the method, ex: u sub

OpenStudy (sshayer):

\[put ~arc \tan \frac{ 1.2 }{ x }=\theta \] \[\frac{ 1.2 }{ x }=\tan \theta ,x=1.2 \cot \theta ,dx=-1.2\csc ^2 \theta~ d \theta \] \[I=\int\limits \cos \theta (-1.2\csc ^2 \theta)d \theta=-1.2\int\limits \frac{ \cos \theta }{ \sin ^2 \theta }d \theta\] \[=-1.2 \int\limits (\sin \theta)^{-2}\cos \theta d \theta=-1.2\frac{ \left( \sin \theta \right)^{-1} }{ -1 }+c\] \[=\frac{ 1.2 }{ \sin \theta }+c\] replace the value of theta

OpenStudy (sshayer):



OpenStudy (issimplcalcus):

would this work if I had a definite integral that I was integrating over distance? Do I just use the derivative and plug in the values?

OpenStudy (issimplcalcus):

not the derivative sorry

OpenStudy (sshayer):

if you are dealing with definite integral ,then find the new limits and no need to come back to x

OpenStudy (issimplcalcus):

Oh thats super smart thanks.

OpenStudy (sshayer):

uw